OpenStudy (anonymous):

movies often use models to create special effects in the godzilla movies the godzilla appears to be 75 feet tall this is achieved by a man in an 8 foot rubber suit photographed in the midst of scale-models of buildings how tall in feet is the model for a building if the building is 50 feet tall in the movie

OpenStudy (anonymous):

i need help

OpenStudy (dumbcow):

set up a proportion.... model size on top, movie size on bottom \[\frac{x}{50} = \frac{8}{75}\] \[x = 8*\frac{50}{75}\] \[x = 8*\frac{2}{3}\] \[x = \frac{16}{3} = 5.33\]

OpenStudy (anonymous):

Thank You so much